Could someone please help me solve:
4^(2-6x)=812

Found 2 solutions by prmarella, funmath:
Answer by prmarella(3)

4^2-6x=812
4*4-6x=812
16-6x=812
-6x=812-16
-6x=796
x=796/(-6)
x= - 398/3

Answer by funmath(2873)

4^(2-6x)=812
ln(4^(2-6x))=ln(812)
(2-6x)ln(4)=ln(812)
(2-6x)ln(4)/ln(4)=(ln(812))/(ln(4))
2-6x=(ln(812)/(ln(4))
-2+2-6x=(ln(812))/(ln(4))-2
-6x=4.832667959-2
-6x=2.832667959
-6x/-6=2.832667959/-6
x=-.4721113264
Happy Calculating!!!!